OpenStudy (anonymous):

Medal Two six-sided dice are rolled. Find the probability of the given outcome (express your answers as fractions in lowest terms). (a) The sum of the dots on the upward faces is greater than 6.

OpenStudy (shaik0124):

total outcomes are 6*6=36 now look for the outcomes where the sum is greater than 6 probability= possible outcomes/total outcome

OpenStudy (shaik0124):

this are the possible outcomes 1,6 2,6 3,6 4,6 5,6 6,6 2,5 3,4 3,5 4,3 4,4 4,5 5,2 5,3 5,4 5,5 6,1 6,2 6,3 6,4 6,5

OpenStudy (shaik0124):

probability=21/36 =7/12

OpenStudy (anonymous):

so the answer is 7/12?

OpenStudy (shaik0124):

1,6 indicates a 1 on the face of first die and 6 on the second die .then the sum will be equal to 1+6 =7 which is greater than 6 so we want all the outcomes which are >6 the number of such outcomes which are greater than 6 is 21

OpenStudy (shaik0124):

understood?

OpenStudy (shaik0124):

yes answer is 7/12 what are the options for ur question

OpenStudy (shaik0124):

reply fast

OpenStudy (anonymous):

i wasnt given any options but for example if it asked for the sum of the dots on the upward faces is greater than 4 it would be 5/6

OpenStudy (shaik0124):

do the same way as i did for >6